a truck hits a brick wall with a force of 100000 kg.m/s2. the truck come to a stop in 0.5. calculate the impulse the truck experinced.
1
Expert's answer
2019-04-02T10:48:17-0400

By the definiton of the impulse, we have:


"J = F \Delta t = 100000 \dfrac{kg \cdot m}{s^2} \cdot 0.5 s = 5 \cdot 10^4 kg \cdot \dfrac{m}{s}."

Answer:"J = 5 \cdot 10^4 kg \cdot \dfrac{m}{s}."